Emilio and Rose sell half of their book collection at a garage sale. They donate 1/5 of the remaining books to a resale shop and keep 40 books. How many books did they have to start with?
Answer by ankor@dixie-net.com(22740)

Emilio and Rose sell half of their book collection at a garage sale. They donate 1/5 of the remaining books to a resale shop and keep 40 books. How many books did they have to start with?
:
Let x = amt of books initially
:
x = yard sale + donated + 40
x = .5x + .2(x -.5x) + 40
x = .5x + .2(.5x)
x = .5x + .10x + 40
x = .6x + 40
x - .6x = 40
.4x = 40
x = 40/.4
x = 100 books
:
:
Check:
.5(100) + .2(50) + 40 =
50 + 10 + 40 = 100